OpenStudy (anonymous):

integral of (3x-1)/(3x+2)

hartnn (hartnn):

@salimal Hi, \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \) For your Question, you can adjust the numerator 3x-1 as 3x-1 = 3x+2 -3 did you get this adjustment ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i get it thank you :)

OpenStudy (amistre64):

.... now hartnn is just showing off

hartnn (hartnn):

welcome ^_^ can you solve further ? lol @amistre64 :P

OpenStudy (anonymous):

hehe yes :P

hartnn (hartnn):

great!

OpenStudy (anonymous):

pellett i couldn`t solve it anyway :/

hartnn (hartnn):

show your steps ? we'll verify them...and spot the error if any

OpenStudy (anonymous):

it`s ok now ..hehe i get it right..just having a bad day :p thank you :)

hartnn (hartnn):

no poblem :)

OpenStudy (anonymous):

Alternatively, you can make the substitution \[u=3x+2\Rightarrow 3x=u-2\\ \frac{1}{3}du=dx\] \[\int\frac{3x-1}{3x+2}dx=\frac{1}{3}\int\frac{u-2-1}{u}du\] Then, integrate term-by-term.

hartnn (hartnn):

^yes, don't forget to re substitute back u =3x+2